how many wavelengths of Kr{atomic number 86}are there in one metre

Grade:

2 Answers

VENKATESH
6 years ago
 Wavelength of Kr 86 is known to be 605 nm (orange in visible region )

So number of wavelengths in 1 meter = 1/605 * 109

Number of Wavelengths = 1652892.562 

adrija biswas
6 years ago
hf=10.2(z-1)^2 …..find f(frequency) then f=c/wavelength get wavelength then divide from 1 metre …...h=plancks const   z=atomic no. c=speed of light in vacuum